OpenStudy (anonymous):

Find teh derivative 4th root of[(1+tanh x)/ (1-tanhx)]

OpenStudy (anonymous):

\[\sqrt[4]{(1+\tanh x) / (1 - \tanh x)}\]

OpenStudy (anonymous):

find the derivative of taht expression and simplify where possible

OpenStudy (anonymous):

\[\sqrt[4]{\frac{1+\text{Tanh}[x]}{1-\text{Tanh}[x]}}=e^{x/2} \]The derivative of the above expression is:\[\frac{e^{x/2}}{2} \]

OpenStudy (anonymous):

how did u get taht